SYLLABUS FOR MATH 416, SPRING 2010

Tuesday, Thursday 11:00 a.m. - 12:15 p.m. MTH 0106

Prerequisites: MATH 141, MATH 240, or permission of the instructor. Familiarity with MATLAB is also required.

Text: Mathematics for Multimedia, Mladen Victor Wickerhauser, Birkhauser - Springer, 2009 (Required)

Instructor: Dr. Wojciech Czaja Office: MTH 4406 Office Hours: Tuesday, Thursday 10:00 a.m. - 10:50 a.m., Wednesday 1:00 - 2:00 p.m., or by appointment Phone: (301) 405 - 5106 Email: wojtek (at) math (dot) umd (dot) edu Class Web Page: www.math.umd.edu/∼wojtek/416 10.htm

Exams: There will be 3 midterm exams during the semester, each for the total of 50 points. The lowest score is dropped. They are tentatively scheduled on February 25, April 1, and May 4. The final exam is scheduled for May 13, 8 a.m. - 10 a.m. It is worth 100 points.

Calculators: Calculators will NOT be allowed for the exams, but small (5 x 3 in.) note cards will be allowed during the exams.

Homework: Every other Friday of the week that has no exam scheduled, homework will be assigned. (Homeworks are due the next Wednesday atthe start of the class.) Each one is worth 20 points, and there will 6 such assignments. The lowest score (including absences) is dropped. Late homework will not be accepted.

Grading: The maximum point total is 300 points and the total used to calculate the final grade is the minimum of the two numbers: 300 points and the largest score in the class. The setting of letter grades will be based on the number of points and will be no worse than: 50% - D, 60% - C, 75% - B, 90% - A.

Academic integrity: The University makes me remind you about its academic integrity policies. So I do. Nobody, however, has to remind me that part of my job is to make sure these policies are obeyed.

Attendance and absences: You are responsible for the material covered in class, whether you attend or not. You are also responsible for the announcements made during class; they may include changes in the syllabus. The instructor will adhere strictly to the official university policy on makeup exams. In particular, it is the policy of the University to excuse the absences of students that result from the following causes: illness of the student, or illness of a dependent; religious observance (where the nature of the observance prevents the student from being present during the class period); participation in university activities at the request of University authorities; and compelling circumstance beyond the student’s control. Students claiming excused absence must apply in writing and furnish documentary support for their assertion that absence resulted from one of these causes. Moreover, foreseeable absences (such as those resulting from religious holidays or participation in university-sponsored events) must be submitted in writing to the instructor by Feb. 15.

Disabilities: If you have a disability disallowing you to test under the usual time or in-class conditions, you may contact the office of Disabled Students Services (DSS) in Shoemaker. (Please let me know by Feb. 15 if you think you may require these services.) If they assess you as meriting private conditions and/or extra time, then you may arrange to take your tests at DSS, with extra time as they indicate. You must arrange this well in advance of a test (in particular: no retakes).

Emergency closures: In case of an emergency that closes the University for an extended period of time (for example, due to inclement weather), be sure to access your email for instructions from me. Also check the University’s home page or call 301-405-SNOW for snow closure information. Should any classes or exams be cancelled, please check the class schedule page for updated schedule information.

Course evaluations: Your participation in the evaluation of courses through CourseEvalUM is important to us, and helps improve teaching and learning at the University. CourseEvalUM will be open for you to complete your evaluations for the Spring semester courses between Tuesday, April 27, and Wednesday, May 12. Please provide feedback on the course, the professor, and the TA before noon on Sunday, December 13, 2009. Evaluations are anonymous and will not be available to faculty and TAs until next semester, so they cannot possibly affect your grade.

You can go directly to the website (www.courseevalum.umd.edu) to complete your evaluations starting December 1. By completing all of your evaluations each semester, you will have the privilege of accessing the summary reports for thousands of courses online at Testudo.

OUTLINE OF MATERIAL

Background Material: Numbers and Computer Arithmetic; Vector Spaces and Linear Transformations [2 weeks]

Fourier Series, Discrete Fourier Transform [4 weeks]

Sampling, Quantization, Precision, and Accuracy [2 weeks]

Wavelet Bases, Discrete Wavelet Transform, [4 weeks]

Applications to Coding, Detection, and Compression [3 weeks]